OpenStudy (anonymous):

fan and trophy Which is an example of an algebraic expression? A. g > 14 B. x + 4 = 13 C. m – 37 < 52 D. b + 32

OpenStudy (anonymous):

@katkipe

OpenStudy (anonymous):

Hmmmm

OpenStudy (anonymous):

Give me a sec

OpenStudy (anonymous):

I belive it is.......... B!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but lets see what @katkipe thinks

OpenStudy (anonymous):

its d

OpenStudy (anonymous):

i know what your doing i did that its D trust me

OpenStudy (anonymous):

Thanks for the medal and fan! :D

OpenStudy (anonymous):

D

OpenStudy (anonymous):

was it B

OpenStudy (anonymous):

for sure @Caitlynsmith6427

OpenStudy (anonymous):

I'd go with B too

OpenStudy (anonymous):

idk yet i have to finish @Caitlynsmith6427

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

do you do k12 6th garde @studygirl12345

OpenStudy (anonymous):

i do k12 7th grade

OpenStudy (anonymous):

ok well i do k12 6th and i had to do that lesson like yesterday

OpenStudy (anonymous):

really that is so weird lol @Caitlynsmith6427

OpenStudy (anonymous):

yep haha i seen it i was like hey i did that question yesterday hah

OpenStudy (anonymous):

may b u could help me then

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

so is d the answer

OpenStudy (anonymous):

i think so im pretty sure

OpenStudy (anonymous):

wait r u doin school right now

OpenStudy (anonymous):

umm i need help with a math problem haha this one Which measure is equivalent to 48 ft? A. 4 in. B. 16 in. C. 288 in. D. 576 in.

OpenStudy (anonymous):

4 @Caitlynsmith6427

OpenStudy (anonymous):

its not 4

OpenStudy (anonymous):

oh ok thanks >.<

OpenStudy (anonymous):

oh thats weird i did 12+12 is 24 then 24+24=48 @Caitlynsmith6427

OpenStudy (anonymous):

Which statements are true according to the order of operations? Choose exactly two answers that are correct. A. (20 – 5) • 3 = 15 • 3 = 45 B. 4 + 12 ÷ 2 = 4 + 6 = 10 C. 15 • (6 ÷ 3) = 90 ÷ 3 = 30 D. 32 + 8 ÷ 4 = 40 ÷ 4 = 10

OpenStudy (anonymous):

umm........ i cant see it haha >.<

OpenStudy (anonymous):

wait can u c all the ? marks

OpenStudy (anonymous):

its acting up on my computer

OpenStudy (anonymous):

no hold on let me work it out

OpenStudy (anonymous):

Choose exactly two answers that are correct. 3. Which statements are true according to the order of operations? Choose exactly two answers that are correct. A. (20 – 5) • 3 = 15 • 3 = 45 B. 4 + 12 ÷ 2 = 4 + 6 = 10 C. 15 • (6 ÷ 3) = 90 ÷ 3 = 30 D. 32 + 8 ÷ 4 = 40 ÷ 4 = 10

OpenStudy (anonymous):

that is the same 1

OpenStudy (anonymous):

its d and B

OpenStudy (anonymous):

Evaluate. (5 • 8 – 12)(36 ÷ 9 • 4)

OpenStudy (anonymous):

its has th ? marks

OpenStudy (anonymous):

so just wrie the problem out

OpenStudy (anonymous):

Sorry, I wasnt online. I think its D

OpenStudy (anonymous):

w

OpenStudy (anonymous):

r

OpenStudy (anonymous):

i

OpenStudy (anonymous):

t

OpenStudy (anonymous):

e

OpenStudy (anonymous):

i

OpenStudy (anonymous):

t

OpenStudy (anonymous):

i tryed lol

OpenStudy (anonymous):

l

OpenStudy (anonymous):

i

OpenStudy (anonymous):

k

OpenStudy (anonymous):

e

OpenStudy (anonymous):

that

OpenStudy (anonymous):

I cant see like anything, I see lots of question marks.

OpenStudy (anonymous):

(5*8-12)(36 divided by 9*4) evaluate

OpenStudy (anonymous):

is that better

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

Okay, try it out..cause its kinda easy.

OpenStudy (anonymous):

my computer says 448

OpenStudy (anonymous):

but i think its 16

OpenStudy (anonymous):

No wait I think it's 608

OpenStudy (anonymous):

yeah i really dont know hah and im in a lower grade so she might be able to do it shes higher haha

OpenStudy (anonymous):

okay parenthases first so do your first set of parenthasis to get 28. then multiply in the next set which will give you 36 and 36 divided by 36= what?

OpenStudy (anonymous):

@Caitlynsmith6427

OpenStudy (anonymous):

ok the answer is 488 P.S. use the order of operation...

OpenStudy (anonymous):

i said 448